### v1

@@ Abstract
1 Background
Understanding the evolution of user behavior through complex interactions in online social networks has been an important topic in computational social science [21, 17, 7, 9, 3, 11], and computer science [28, 16, 22, 26]. However, researchers face two challenges in studying user behavior on social media platforms: (1) platforms typically design recommendations to target users, making it difficult to isolate their effects on user behavior; and (2) platforms often limit access to detailed timing and comprehensive records of how users interact. Table 1 categorizes and compares how eight social media platforms manage key aspects of users’ social interactions, including following (establishing one-way connections and receiving updates from connected users), blocking (preventing visibility and engagements between users), creating (generating a new community), and joining (subscribing and receiving a community’s suggested content).
Table 1: Comparing accessible social interactions of users on social media platforms via APIs. The symbols used are: $\times$ for complete exclusion, $\circ$ for imperfect coverage (i.e., no timestamp), and $\bullet$ for perfect coverage.
@@ 1 Background
Twitter [27, 19] $\circ$ $\times$ $\times$ $\times$
@@
LinkedIn [6, 30] $\times$ $\times$ $\times$ $\times$
Reddit [2, 10] $\times$ $\times$ $\circ$ $\circ$
@@
Bluesky Social $\bullet$ $\bullet$ $\bullet$ $\bullet$
Interest in alternatives to the central control of traditional social media platforms has led to a substantial migration of users to decentralized platforms [14]. Here, we focus on Bluesky [^1] https://bsky.social/about, a social media platform currently boasting over 5 million active users, where data are distributed and managed across multiple providers. Despite its decentralized architecture, Bluesky offers a seamless user experience comparable to that of centralized platforms. This is achieved through interoperability facilitated by its protocol, ATproto [18]. Bluesky features a default reverse chronological timeline and allows users to add customized recommendation algorithms created by other users. This user-driven design facilitates the study of user behavior without the potential control exerted by platforms [25].
@@
2 PLATFORM INTERFACE
To gain a clearer understanding of Bluesky’s user experience, Figure 2 illustrates social interactions using the corresponding authors’ account as an example. On the right side of the figure, the follow action represents a user’s interest in another user, allowing them to receive updates and posts in their default timeline. The block feature enhances privacy and safety by ensuring that users who have blocked each other cannot view or engage with each other’s posts, making them invisible to both sides. On the left side of the figure, there is a list of Feeds, which are dynamic content streams with customizable algorithms. Users join Feeds of interest, like philsky (Philosophy content) and Discover (Trending content), by clicking the heart icon below their display names and descriptions, an action known as Feed like. Each Feed is created by a user using ATproto’s Feed generator.
3 COLLECTION METHODS AND DESIGN
We used the Bluesky API provided by ATproto [^4] https://atproto.com/ protocol. As shown in Figure 3, there are three sequential steps in the data collection process: (1) collecting existing Feeds from Bluesky at the community level; (2) gathering information about users who generated or liked these Feeds at the user-to-community level; and (3) compiling relationships between users by examining if they blocked or followed each other at the user-to-user level.
4 RECORDS AND STORAGE
4.1 Storage Principles
In accordance with FAIR (Findability, Accessibility, Interoperability, and Reuse of digital assets) principles [8], we took measures to ensure that the data we utilized or curated are devoid of personally identifiable information. Any text-level information about users’ profiles or their posts is not covered in the dataset for privacy measures. Throughout our analysis, we anonymized user identifiers with a MongoDB database through field-level encryption.
4.2 Scope of Records
@@
Figure 3: An illustration of our data collection pipeline on Bluesky is presented in three levels. Each stage of collection captures the timestamp of user behavior with millisecond-level precision in UTC format.
4.3 Definition of User
@@
Table 2 presents statistics of Bluesky, including the number of users (Feed creators and members), the total Feeds, and the engagement counts from Feed members. In order to integrate their connections into a multi-graph through their shared Feeds, we only focused on the interactions among the Feed creators and Feed members
Table 2: Summary of the user and Feed samples identified in Bluesky. Active users and Feeds are set at the time of data collection (May 11, 2024).
Types Number
Active Users 5,659,340
Active Feeds 39,968
Feed Creators (unique active users who have generated at least one active Feed) 17,617
Feed Members (unique active users who have liked at least one active Feed) 147,577
Total Like Count on Feeds by Feed Members 322,930
Total Block Count of Feed Members 2,677,622
Total Follow Count of Feed Members 24,482,447
4.4 Dimensions of Records
Figure 3 shows our data collection process on Bluesky. Based on users, Feeds, and their follow and block relationships at each timestamp of social interaction, we defines three dimensions of user and community interactions.
@@ 4.4 Dimensions of Records
    Feed Creator Interaction Network: A signed graph representing follows (positive edges) and blocks (negative edges) among Feed creators, denoted as $G_{\mathcal{C}}=\{\mathcal{C},\mathcal{E}^{+},\mathcal{E}^{-}\}$, where $\mathcal{C}$ is the set of Feed creators.
@@
    Feed Member Interaction Network: A signed graph representing follows (positive edges) and blocks (negative edges) among users, denoted as $G_{\mathcal{M}}=\{\mathcal{M},\mathcal{E}^{+},\mathcal{E}^{-}\}$, where $\mathcal{M}$ is the set of Feed members.
@@
    Community Interaction Network: An affiliation graph connecting communities of users with Feeds, denoted as $G_{\mathcal{A}}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{\mathcal{C},\mathcal{M},\mathcal{F}\}$ represents three types of nodes (creators, members, and Feeds), and =$\mathcal{E}=\{\mathcal{G},\mathcal{L}\}$ represents two types of undirected edges (Feed generation and Feed likes).
Following and blocking are not mutually exclusive actions. A user can follow others initially and then block them later. For these cases, we prioritize the blocking interaction by retaining the negative edges (blocking) and removing the positive edges (following). Based on this, we present the statistics for the three dimensions of the dataset: (1) Table 3 shows the Feed creator interaction network. The graph contains 349 isolated nodes and 485 node pairs with both negative and positive edges. (2) Table 4 presents the Feed member interaction network. The graph includes 34 isolated nodes and 5,502 node pairs with both negative and positive edges. (3) Table 5 describes the community interaction network. In the affiliation graph, we identified 8,820 users who act as both Feed creators and Feed members. These users are represented as two nodes: one as a member node and the other as a creator node.
Table 3: First dimension – statistics of the signed graph $G_{\mathcal{C}}$, showing blocking and following among the Feed creators identified in Table 2.
Types Number
Node (Feed Creators) 17,617
Negative Edge (Blocking among Feed Creators) 26,992
Positive Edge (Following among Feed Creators) 292,331
Table 4: Second dimension – statistics of the signed graph $G_{\mathcal{M}}$, showing blocking and following among the Feed members identified in Table 2.
Types Number
Node (Feed Members) 147,577
Negative Edge (Blocking among Feed Members) 532,614
Positive Edge (Following among Feed Members) 5,766,570
@@
Types Number
Member Node (Feed Members) 147,577
Creator Node (Feed Creators) 17,617
Feed Node (Active Feeds) 39,968
Join Edge (Liking Feeds) 322,930
Create Edge (Genearting Feeds) 39,968
5 INSIGHTS AND NOTES
5.1 Feed Distribution
Figure 4 shows the popular Feeds and the frequent terms used in the display name of Feeds. As the 64-character limit for Feed names will encourage users to choose words carefully, we assume users selectively choose words in display names to represent the interest of the Feed. By removing redundant words such as feed, tab, post, and test, our analysis revealed that a large number of Feeds are found to focus on topics such as art, media, music, japanese, ai, and nsfw. This indicates that Bluesky users have a focused preference for niche content.
Figure 4: Bar charts showing (A) top 20 most liked Feed’s display names and (B) top 20 most frequently used terms in Feed’s display names.
The popularity of Feed does not necessarily correlate with its prevalence in Bluesky. To examine the relationship between display names and the popularity of Feeds, we calculated term frequency vectors and computed Pearson correlation coefficients for each word. Figure 5 shows words with the highest and lowest correlations of words. Feeds with words like likes, discover, and hot were strongly associated with popularity. Community and social topics such as blacksky (Black community), ukraine (Ukraine-Russian war), academic (science, astronomy, etc), transqueer (LGBTQ), and adult content also showed positive correlations. likes was the highest term, indicating that liking is perceived as an important behavior among users for discovering information on Bluesky. Terms like art, gallery, and media were among the least popular despite their prevalence in Feeds’ display names. However, the correlation for these terms is very low (below $-0.005$), implying minimal impact on Feed’s popularity.
@@ 5.1 Feed Distribution
Summary (Communities) There is a notable contrast between the Feeds users prefer to create and those that achieve widespread popularity. On Bluesky, niche Feeds related to art and media are prevalent. However, Feeds that facilitate content discovery and highlight trending topics tend to garner the most number of likes from users.
Figure 5: Bar charts showing Pearson correlations between term frequency in Feed’s display names and the number of likes: (A) terms with top 20 correlations, (B) terms with bottom 20 correlations.
5.2 Temporal Dynamics
Figure 6 shows the temporal trends of three dimensions of social interactions in our dataset. We identified a sudden surge in both follows and blocks after Bluesky transitioned to an invitation-free platform on February 6, 2024. Among Feed creators, follows increased by 19.2% and blocks by 27.3%. Similarly, among Feed members, follows increased by 15.8% and blocks by 25.9%. After the transition, 44% of the Feeds were generated, and users liked 34.12% of them. We speculate that the larger increase in blocks compared to follows indicates a heightened need for users to filter out potentially unwanted interactions from new users and Feeds, especially after the platform became completely open.
Figure 7 illustrates the evolution of the degree distribution across three dimensions of our dataset. The average and standard deviation of in-degrees and out-degrees gradually increase over time, showing stabilization as time progresses. Despite the surge in users and interactions during the transition to an invitation-free platform, the average and standard deviation of interactions remain stable across all three dimensions. This indicates that Bluesky managed the influx of users and their interactions, maintaining stability and aligning with existing trends during its establishment.
Summary (Temporal Dynamics) After Bluesky’s official public launch, social interactions surged across all dimensions. The rate of user blocks increased more sharply compared to the rate of follows, and the creation of new Feeds outpaced the rate at which users joined them. Despite this surge in users and their interaction levels, Bluesky managed to stabilize the interactions to some extent before the transition to an invitation-free platform.
Figure 6: Temporal trends in the number of nodes (orange) and edges (blue) for three networks: $G_{\mathcal{C}}$ (Feed Creator Interaction), $G_{\mathcal{M}}$ (Feed Member Interaction), and $G_{\mathcal{A}}$ (Community Interaction). The red dotted line marks when Bluesky transitioned to an invitation-free platform.
Figure 7: Temporal trends in degree distributions for three networks: $G_{\mathcal{C}}$ (Feed Creator Interaction), $G_{\mathcal{M}}$ (Feed Member Interaction), and $G_{\mathcal{A}}$ (Community Interaction). Lines show average in-degree (blue) and out-degree (orange); shaded areas indicate standard deviation. The red dotted line marks when Bluesky transitioned to an invitation-free platform.
5.3 Network Analysis
@@
We base our network analysis on the static graph from the last data collection on May 11, 2024. For simplicity, we focus on the largest component of each graph, encompassing 88% of nodes in $G_{\mathcal{C}}$, 94% of nodes in $G_{\mathcal{M}}$, and 86% of nodes in $G_{\mathcal{A}}$. To comprehensively view each network’s topology, we carefully selected applicable measures for the signed graph [4] and measures for the affiliation graph [24].
5.3.1 User-to-User (singed graph) Networks Analysis
@@
Table 6 summarizes metrics for the signed graphs $G_{\mathcal{C}}$ and $G_{\mathcal{M}}$, highlighting five key aspects: (1) Positive assortativity is slightly negative in both graphs, indicating high-degree users tend to follow lower-degree users. (2) Negative degree assortativity is higher for $G_{\mathcal{C}}$ compared to $G_{\mathcal{M}}$, showing distinct blocking patterns between creators and members. (3) Positive reciprocity is higher than negative reciprocity in both graphs, suggesting stronger mutual following than blocking. (4) Both graphs exhibit similar positive clustering coefficients but low negative clustering coefficients, indicating that followers’ neighbors are more likely to be connected than blockers’ neighbors. (5) Structural balancedness is slightly higher in $G_{\mathcal{C}}$ than in $G_{\mathcal{M}}$, suggesting a greater prevalence of social balance, where an enemy’s enemy is considered a friend, among Feed creators.
@@
Diameter The greatest distance between any pair of nodes 12 11
Average Path Length The average shortest path between all pairs of nodes 3.18 3.19
Degree Density Proportion of actual edges compared to possible edges 1.07 $e^{-3}$ 2.97 $e^{-4}$
Positive Degree Assortativity Bias for nodes to follow neighbors with similar following degrees -0.13 -0.03
Negative Degree Assortativity Bias for nodes to block neighbors with similar blocking degrees 0.02 -0.15
Positive Reciprocity Likelihood of nodes to be mutually linked by following edges 0.46 0.50
@@
Positive Clustering Coefficient Average of two adjacent nodes of a follower also follow each other 0.24 0.25
Negative Clustering Coefficient Average of two adjacent nodes of a blocker also block each other 0.03 0.03
Structural Balancedness Ratio of socially balanced triangles based on undirected edges 0.91 0.89
@@
Diameter The greatest distance between any pair of nodes 24 11
Average Path Length The average shortest path between all pairs of nodes 60.49 3.06
Degree Density Proportion of actual edges compared to possible edges 1.87 $e^{-5}$ 1.07 $e^{-2}$
Degree Assortativity Nodes to connect with similar degrees of neighborhood -0.16 0.50
Clustering Coefficient Average of two neighbors of a node are also neighbors to each other 0.80 0.92
5.3.2 User-to-Community (affiliation graph) Networks Analysis
@@
where $U(f_{i})$ denotes the set of any users adjacent to one of Feeds $f_{i}\in\mathcal{F}$ from $G_{\mathcal{A}}$. Two nodes $f_{i}$ and $f_{j}$ are linked if they share at least one user. The weight of an edge $w(f_{i},f_{j})$ is the number of shared users between the Feeds.
@@
Figure 8 illustrates the community analysis of the projected line graph $P_{\mathcal{A}}$, which are clustered using the Leiden Algorithm [29], with annotations derived from Feeds’ descriptions. The clusters reveal a variety of groups representing different languages and regions, such as German, Japanese, Spanish, and English groups on Bluesky. Additionally, the network features clusters centered on political topics, like the Ukraine-Russia war, alongside more casual feeds, with a notable prevalence of art-related communities. As discussed during network analysis, the relationships between Feeds exhibit small-world properties. Further exploration could uncover tightly knit community groups, including higher-order analyses to identify strongly connected clusters and examine their characteristics with respect to language, political geography, and relevant social events.
Summary (Network Analysis) Followings for both creators and members exhibit negative degree assortativity, connecting to nodes with lower degrees. Members also tend to make blockers more connected, with higher structural balancedness highlighting distinct user roles. In community analysis, the projection of the affiliation graph reveals small-world properties, particularly with prominent art-related communities. Studying higher-order relations using the projected line graph appears promising for understanding how the platform maintains its ecosystem between communities.

### v2

@@ Abstract
BACKGROUND
Understanding the evolution of user behavior through complex interactions in online social networks has been an important topic in computational social science [4, 18, 12], and computer science [29, 17, 23]. However, researchers face two challenges in studying user behavior on social media platforms: (1) platforms typically design recommendations to target users, making it difficult to isolate their effects on user behavior; and (2) platforms often limit access to detailed timing and comprehensive records of how users interact. Table 1 categorizes and compares how eight social media platforms manage key aspects of users’ social interactions, including following (establishing one-way connections and receiving updates from connected users), blocking (preventing visibility and engagements between users), creating (generating a new community), and joining (subscribing to a community’s suggested content).
Table 1: Comparing accessible social interactions of users on social media platforms via APIs. The symbols used are: $\times$ for complete exclusion, $\circ$ for partial access without timestamps, and $\bullet$ for perfect coverage.
@@
Twitter/X [27, 21] $\circ$ $\times$ $\times$ $\times$
@@
Threads [15] $\times$ $\times$ $\times$ $\times$
LinkedIn [7] $\times$ $\times$ $\times$ $\times$
Reddit [3] $\times$ $\times$ $\circ$ $\circ$
@@
Bluesky [2, 28] $\bullet$ $\bullet$ $\bullet$ $\bullet$
Interest in alternatives to the central control of traditional social media platforms has led to a substantial migration of users to decentralized platforms [15, 9]. Here, we focus on Bluesky [^1] https://bsky.social/about, a social media platform currently boasting over 5 million active users, where data are distributed and managed across multiple providers. Despite its decentralized architecture, Bluesky offers a seamless user experience comparable to that of centralized platforms. This is achieved through interoperability facilitated by its protocol, ATproto [19]. Bluesky features a default reverse chronological timeline and allows users to add customized recommendation algorithms created by other users. This user-driven design facilitates the study of user behavior on social media without the potential control exerted by platforms [26].
@@
PLATFORM INTERFACE
Figure 2 illustrates social interactions using the corresponding authors’ account as an example. On the right side of the figure, the follow action represents a user’s interest in another user, allowing them to receive updates and posts in their default timeline. The block feature enhances privacy and safety by ensuring that users who have blocked each other cannot view or engage with each other’s posts, making them invisible to both sides. On the left side of the figure, there is a list of Feeds, which are dynamic content streams with customizable algorithms. Users create and join Feeds, like Philosophy (Philosophy content) and Discover (Trending content). A user creates each feed through ATproto’s Feed Generator and joins Feeds of interest by clicking the heart icon on the feed’s profile, an action called Feed Like.
COLLECTION METHODS AND DESIGN
We used the Bluesky API provided by ATproto [^4] https://atproto.com/ protocol. As shown in Figure 3, there are three sequential steps in the data collection process: (1) collecting existing Feeds from Bluesky at the community level; (2) gathering information about users who generated or liked these Feeds at the user-to-community level; and (3) compiling relationships between users by examining if they blocked or followed each other at the user-to-user level. We ensure data quality and retrieval efficiency by using one of REST API endpoints named as public.api.bsky.app/xrpc/com.atproto.repo.listRecords, which is publicly accessible without authentication. The records are managed within Bluesky Social’s PDS named bsky.social.
RECORDS AND STORAGE
STORAGE PRINCIPLES
We implemented privacy measures, including field-level encryption to anonymize user IDs and the removal of personally identifiable information such as usernames, display names, and biographies, to protect user privacy. At the same time, we adhered to the FAIR [10] principles (Findability, Accessibility, Interoperability, and Reusability) by ensuring that the dataset remains reusable and well-documented with appropriate metadata, making it both privacy-compliant and suitable for future research. A decryption key for anonymized user IDs will be made available to bona fide researchers upon request.
SCOPE OF RECORDS
@@
Figure 3: An illustration of our data collection pipeline for BlueTempNet is shown in three levels. Each stage of collection captures the timestamp of a user’s interactions with millisecond-level precision in UTC format.
USER CATEGORIES
@@
Table 2 provides basic statistics for Bluesky, including the number of users (comprising Feed creators and members), the number of Feeds, and some engagement metrics associated with their profiles. It is important to note that these numbers do not represent our collected dataset. Subsequent section details our specific method on collecting the interactions between Feed creators and members.
Table 2: Overview of total users and Feeds on Bluesky as of May 11, 2024. Profile metrics, such as #Likes, are aggregated to show overall totals.
Types Value
#Users 5,659,340
#Feeds 39,968
#Likes on Feeds 322,930
#Feed Creators (unique users who have generated at least one Feed) 17,617
#Feed Members (unique users who have liked at least one Feed) 147,577
DIMENSIONS OF RECORDS
Figure 3 illustrates the data collection process of BlueTempNet. Based on the user types we categorized, we propose three network dimensions to elicit the hierarchical structure of our dataset and their various interactions.
@@
    Feed Creator Interaction Network: This is a signed, directed graph representing interactions between Feed creators. It is defined as $G_{\mathcal{C}}=\{\mathcal{C},\mathcal{E}^{+},\mathcal{E}^{-}\}$, where $\mathcal{C}$ denotes the set of Feed creators, $\mathcal{E}^{+}$ represents the set of positive edges (following), and $\mathcal{E}^{-}$ represents the set of negative edges (blocking).
@@
    Feed Member Interaction Network: This is a signed, directed graph representing interactions between Feed members. It is defined as $G_{\mathcal{M}}=\{\mathcal{M},\mathcal{E}^{+},\mathcal{E}^{-}\}$, where $\mathcal{M}$ denotes the set of Feed members, $\mathcal{E}^{+}$ represents the set of positive edges (following), and $\mathcal{E}^{-}$ represents the set of negative edges (blocking).
@@
    Community Interaction Network: This is affiliation graph that connects communities of users to Feeds. It is defined as $G_{\mathcal{A}}=\{\mathcal{V},\mathcal{E}\}$, where the set of nodes $\mathcal{V}=\{\mathcal{C},\mathcal{M},\mathcal{F}\}$ consists of three distinct types: Feed creators, Feed members, and Feeds. The set of edges =$\mathcal{E}=\{\mathcal{G},\mathcal{L}\}$ consists of two undirected edge types: generation, denoted by $\mathcal{G}$, and likes, denoted by $\mathcal{L}$.
Following and blocking are not mutually exclusive actions. A user can follow others initially and then block them later. For these cases, we prioritize the blocking interaction by retaining the negative edges (blocking) and removing the positive edges (following). Based on this, we present the statistics for the three dimensions of the dataset: (1) Table 3 shows the Feed creator interaction network. The graph contains $1,749$ isolated nodes and there were $485$ edges with both negative and positive signs. (2) Table 4 presents the Feed member interaction network. The graph includes $9,145$ isolated nodes and there were $5,502$ edges with both negative and positive signs. (3) Table 5 describes the community interaction network. In the affiliation graph, we identified $8,527$ users who act as both Feed creators and Feed members. These users are represented as two nodes: one as a member node and the other as a creator node.
Figure 4: Scatter plot displaying the term frequency of terms in Feeds display names (y-axis) and their Pearson correlations with the number of likes received (x-axis). Scatter points are color-coded based on the direction of the correlation. We tagged the points for the top 10 highest and lowest Pearson correlations by displaying the corresponding terms.
Table 3: First dimension – statistics of the signed directed graph $G_{\mathcal{C}}$, showing block and follow among the Feed creators identified in Table 2.
Types Count
Node (Feed Creators accessible through API) 17,146
Negative Edge (Blocking among Feed Creators) 24,362
Positive Edge (Following among Feed Creators) 273,696
Table 4: Second dimension – statistics of the signed directed graph $G_{\mathcal{M}}$, showing block and follow among the Feed members identified in Table 2.
Types Count
Node (Feed Members accessible through API) 134,946
Negative Edge (Blocking among Feed Members) 435,700
Positive Edge (Following among Feed Members) 4,871,132
@@
Types Count
Creator Node (Feed Creators) 17,536
Member Node (Feed Members) 136,724
Feed Node (Feeds accessible through API) 39,235
Join Edge (Users Liking Feeds) 297,828
Create Edge (Users Generating Feeds) 39,235
INSIGHTS AND NOTES
FEED ECONOMIC DYNAMICS
Figure 4 illustrates the most common terms used in Feed display names and their correlation with popularity, measured by likes. This analysis used term frequency (TF) vectors, defined as $\text{TF}(t,d)=\frac{\text{count}(t,d)}{\sum_{t^{\prime}\in d}\text{count}(t^{\prime},d)}$, where $t$ is the term and $d$ is the display name. Pearson correlation coefficients were calculated between the occurrence of these terms and the number of likes received by Feeds containing them. Interestingly, terms such as art, media, and gallery, though commonly used in display names, showed little to no correlation with popularity, with most correlations falling below $-0.005$, suggesting a negligible association with Feed popularity. In contrast, terms like likes, discover, classic, and hot exhibited strong positive correlations with Feed popularity. In addition, we found that terms associated with specific communities, such as blacksky (Black community), science (academic communication), booksky (book club), and adult content, were also positively correlated with higher numbers of likes.
@@
Summary (Feed Economic Dynamics) There is a contrast between the niche Feeds users create and those that gain widespread popularity. On Bluesky, while art and media Feeds are common, those centered on content discovery, trending topics, and certain communities often receive the most likes.
TEMPORAL INTERACTION DYNAMICS
Figure 5 highlights the trends in three dimensions of social interactions in BlueTempNet after the platform’s shift to an invitation-free model on February 6, 2024. From February 6 to the conclusion of our analysis, the cumulative distribution shows that the number of follows increased by $19.2\%$ and blocks by $27.3\%$ in the first dimension. In the second dimension, follows rose by $15.8\%$ and blocks by $25.9\%$. The third dimension indicates a remarkable $44\%$ increase in the creation of Feeds, along with $34.12\%$ rise in joining Feeds through likes. This analysis demonstrates that blocking users and creating Feeds became more pronounced than their counterpart during the the platform’s public launch.
Figure 6 shows how the distribution of social interactions evolves over time, based on the coefficient of variation (CV) to measure distributional stability. The CV, defined as the ratio of the standard deviation to the mean, increases steadily across all interactions, with a noticeable rise after October 2023. A significant shift follows Bluesky’s transition to open beta in February 2024. After the public launch, follow and block interactions remain relatively stable, while create and join interactions show increased variability before stabilizing. This contrast reveals that user-to-user interactions are more resilient than user-to-community interactions during periods of significant user growth.
Summary (Temporal Interaction Dynamics) After Bluesky’s public launch, social interactions among users surged across all dimensions. Blocks increased at a faster rate than follows, and the creation of Feeds outpaced the rate of users joining Feeds. The coefficient of variation indicated stable growth patterns in follow and block interactions, while create and join interactions showed greater variability. This underscores the distinct characteristics of user behaviors during the platform’s grand opening.
Figure 5: Trends in the increase of new edges across three networks: $G_{\mathcal{C}}$, $G_{\mathcal{M}}$, and $G_{\mathcal{A}}$. The lines represent the number of new edges added on each date, while the shaded area indicates the cumulative distribution function (CDF) for each interaction type. The red dotted line marks Bluesky’s transition to an invitation-free platform.
Figure 6: Trends of the coefficient of variation (CV) based on degrees of nodes across three networks: $G_{\mathcal{C}}$, $G_{\mathcal{M}}$, and $G_{\mathcal{A}}$. Solid and dashed lines indicate in-degree and out-degree, respectively. The red dotted line marks Bluesky’s transition to an invitation-free platform.
NETWORK ANALYSIS
@@
We base our network analysis on the static graph from the last data collection on May 11, 2024. For simplicity, we focus on the largest component of each graph, encompassing $89\%$ of nodes in $G_{\mathcal{C}}$, $92\%$ of nodes in $G_{\mathcal{M}}$, and $86\%$ of nodes in $G_{\mathcal{A}}$. To comprehensively view each network’s topology, we carefully selected applicable measures for the signed graphs [5] and measures for the affiliation graph [25].
USER-TO-USER (SIGNED GRAPH)
@@
Table 6 summarizes metrics for the signed graphs $G_{\mathcal{C}}$ and $G_{\mathcal{M}}$, highlighting five key aspects: (1) Positive assortativity is slightly negative in both graphs, indicating high-degree users tend to follow lower-degree users. (2) Negative degree assortativity is higher for $G_{\mathcal{C}}$ compared to $G_{\mathcal{M}}$, showing distinct blocking patterns between creators and members. (3) Positive reciprocity is higher than negative reciprocity in both graphs, suggesting stronger mutual following than blocking. (4) Both graphs exhibit similar positive clustering coefficients but low negative clustering coefficients, indicating that followers’ neighbors are more likely to be connected than blockers’ neighbors. (5) $G_{\mathcal{M}}$ shows a slightly higher proportion of balanced triangles than $G_{\mathcal{C}}$, indicating greater prevalence of social balance among Feed members [8, 5].
@@
Diameter The greatest distance between any pair of nodes 12 13
Average Path Length The average shortest path between all pairs of nodes 3.26 3.29
Degree Density Proportion of actual edges compared to possible edges 1.89 $e^{-3}$ 5.10 $e^{-4}$
Positive Degree Assortativity Bias for nodes to follow neighbors with similar following degrees -0.08 -0.06
Negative Degree Assortativity Bias for nodes to block neighbors with similar blocking degrees -0.21 -0.16
Positive Reciprocity Likelihood of nodes to be mutually linked by following edges 0.45 0.51
@@
Positive Clustering Coefficient Average of two adjacent nodes of a follower also follow each other 0.18 0.20
Negative Clustering Coefficient Average of two adjacent nodes of a blocker also block each other 0.04 0.04
Structural Balance of Triangles Ratio of socially balanced triangles based on undirected edges 0.47 0.52
@@
Diameter The greatest distance between any pair of nodes 21 10
Average Path Length The average shortest path between all pairs of nodes 5.30 3.25
Degree Density Proportion of actual edges compared to possible edges 2.29 $e^{-5}$ 1.40 $e^{-2}$
Degree Assortativity Nodes to connect with similar degrees of neighborhood -0.16 0.59
Clustering Coefficient Average of two neighbors of a node are also neighbors to each other 0.00 0.77
USER-TO-COMMUNITY (AFFILIATION GRAPH)
@@
where $U(f_{i})$ denotes the set of users adjacent to one of Feeds $f_{i}\in\mathcal{F}$ from $G_{\mathcal{A}}$. Two nodes $f_{i}$ and $f_{j}$ are linked if they share at least one user. The weight of an edge $w(f_{i},f_{j})$ is the number of shared users between the Feeds.
@@
Figure 7 presents the community analysis of the projected line graph $P_{\mathcal{A}}$, where clusters are identified through the Leiden Algorithm [30] designed for a weighted graph. Each cluster is visualized through word clouds generated from the descriptions in the Feeds, highlighting communities based on different languages and regions, including German, Japanese, Spanish, and English. Furthermore, the clusters features communities for both political and casual topics, such as the Ukraine-Russia war and art-related communities. As discussed during network analysis, this projected line graph exhibits small-world properties. Further exploration could uncover tightly knit community groups, including higher-order analyses to identify strongly connected clusters and examine their characteristics with respect to language, political geography, and relevant social events.
Summary (Network Analysis) Both creators and members exhibit negative degree assortativity in their followings, primarily linking to nodes with lower degrees. Additionally, members often connect blockers more effectively, which highlights distinct user roles and higher structural balance. Projecting the affiliation graph revealed small-world properties within Feeds, particularly clustered around communities related to art and languages.